# Neural Receptive Fields, Stimulus Space Embedding and Effective Geometry of Scale-Free Networks

### v1

@@ VI Methods
We employ the approach of Kryukov et al. [52, 53], which allows one to generate random scale-free graphs $G(V,E)$ with controllable parameters, including the power-law degree distribution exponent $\gamma$. The network nodes are embedded in hyperbolic space $\mathbb{H}^{d}$ (with $d\geq 2$), where each node is assigned one hyperbolic radial coordinate $r$ and d1d-1 angular coordinates $\boldsymbol{\theta}^{\mathrm{emb}}=(\theta^{\mathrm{emb}}_{1},\theta^{\mathrm{emb}}_{2},\dots,\theta^{\mathrm{emb}}_{d-1})$, reflecting an effective geometric structure.
For simplicity, we consider stimulus spaces that define an $d^{\mathrm{stim}}$-dimensional orientation (sphere $\mathbb{S}^{d^{\mathrm{stim}}}$). A stimulus is defined as a scalar field over $\mathbb{S}^{d^{\mathrm{stim}}}$, for example, exhibiting a cosine-like profile in the stimulus coordinates $\boldsymbol{\theta}^{\mathrm{stim}}=(\theta_{1}^{\mathrm{stim}},\theta_{2}^{\mathrm{stim}},\dots,\theta_{d^{\mathrm{stim}}}^{\mathrm{stim}})$. In both rate-based and spiking models of neural activity (see the next subsection), the $d^{\mathrm{stim}}$-dimensional stimulus space is associated with the d1d-1 -dimensional embedding boundary via the mapping $\boldsymbol{\theta}^{\mathrm{emb}}\mapsto\boldsymbol{\theta}^{\mathrm{stim}}$, under the dimensionality-matching condition $d^{\mathrm{stim}}=d-1$.
@@
where $f^{\mathrm{stim}}_{i}(t)$ is the value of the stimulus field in the receptive coordinates of neuron $i$, and $\xi_{i}(t)\sim\mathcal{N}(0,1)$ is Gaussian white noise. The parameters $g^{\mathrm{stim}}=5$, $g^{e,\mathrm{noise}}=1$, and $g^{i,\mathrm{noise}}=4$ scale the contributions of stimulus and noise to excitatory and inhibitory neurons, respectively. As a stimulus field for detecting receptive fields in the spiking model, we use a narrow profile that smoothly decays toward the periphery, defined as $I^{\mathrm{stim}}(\boldsymbol{\theta})=\exp\left(-\|\boldsymbol{\theta}_{\mathrm{stim}}-\boldsymbol{\theta}\|/w_{\mathrm{stim}}\right)$, where $\boldsymbol{\theta}_{\mathrm{stim}}$ denotes the current peak location on the stimulus sphere, and $w_{\mathrm{stim}}$ controls the spatial width of the stimulus.
@@
where $S\in\mathbb{R}^{N\times N}$ is the internal connectivity matrix that determines how the nodes influence each other. Excitatory connections are defined according to the adjacency matrix of a scale-free graph generated and are scaled by excitatory conductance $g^{e}$. All remaining entries in the matrix, that is, those that do not correspond to excitatory connections, represent inhibitory influences and are randomly sampled from a uniform distribution in the interval [1,0][-1,0], then scaled by inhibitory conductance $g^{i}$. In this way, the matrix $S$ simultaneously encodes structured excitatory input and homogeneous random inhibitory connectivity.
Synaptic depression modifies transmission through the probability of release $P_{\mathrm{rel}}(t)$  [69], which evolves as
@@
where $\tau_{\mathrm{depr}}=500$ ms is the recovery time constant, $P_{0}=1$ is the baseline release probability, and $f=0.8$ is the depression factor. The depression factor $f=0.8$ describes how much the synapses depress after each spike, effectively modifying the release probability via $P_{\mathrm{rel}}\rightarrow fP_{\mathrm{rel}}$.
The firing rate is determined by a simple threshold function $r(t)=[I(t)]_{+},$ where $\cdot$$_{+}$ denotes rectification, corresponding to taking the non-negative part: $x]_{+}=\max(x,0)$.
@@
A full description of the methodology is provided in  [65]. The CA1 hippocampal activity was recorded in rats navigating a linear track with a total length of approximately $5$ meters, composed of ten $\sim 50$ cm sections. These segments were arranged into a flexible U-shaped track. The two arms of the track moved independently by remotely operated stepper motors. The displacement varied pseudo-randomly across trials, allowing to test a large set of configurations.
To control distal visual cues, experiments were conducted in complete darkness. To support navigation, the track surface was coated with a scentless, washable, nontoxic glow-in-the-dark paint, which provided a faint green glow. A large screen blocked the view of one glowing arm from the other, ensuring the rat could only see the arm it was traversing. Stepper motors operated quietly, and the animals showed no overt behavioral responses to the associated noise or vibrations during track repositioning.
Each animal was recorded for $7$ days on the track, performing on average about $40$ laps during each run session. The average traversal time ($\sim 34$ s) did not significantly differ from pre-training performance on the static track. More details can be found in [65].
@@
All analyses excluded epochs when animals ran at speeds <<2 cm/s. To compare spiking activity in planar and linear reference frames, we used Cartesian 2D coordinates and then linearized the trajectory such that the $x$ coordinate represented distance from one of the food wells, ignoring displacements orthogonal to the track direction.
Receptive fields were estimated from single-unit spiking activity along the linear track. For each cell, spike positions were binned along the $x$-coordinate and converted into a spatial firing-rate profile by normalizing spike counts with the occupancy, i.e. the time spent by the animal in each spatial position. The resulting firing-rate profiles were smoothed with Gaussian filtering to reduce noise while preserving the structure of place fields. The reception fields were then defined as contiguous regions where the smoothed firing rate exceeded a fixed threshold of 2 Hz above baseline activity. For consistency of analysis, only a single receptive field was retained per neuron: in cases where multiple fields were detected, all but the largest contiguous field containing the peak firing location were discarded. Data analysis was performed in Python.
To quantify the curvature of the receptive field distribution, we followed the exponential model of place-field sizes described by Eq. (1):
@@
where $s$ denotes the field size, $\xi$ is the global curvature parameter, and $s_{\min},s_{\max}$ are the minimal and maximal observed field sizes.
To estimate the local curvature $\xi_{\mathrm{loc}}$ along the track, we applied this formulation within sliding windows (40 cm) on the spatial axis $x$. In each window, receptive field sizes $s$ with centers located inside the window were extracted, and a maximum likelihood estimate of $\xi$ was computed using the truncated exponential model above. This procedure yielded a spatially resolved profile $\xi_{\mathrm{loc}}(x)$, capturing local variations in field-size scaling. The resulting $\xi_{\mathrm{loc}}(x)$ values were subsequently smoothed with a Gaussian kernel to suppress noise while preserving spatial trends.

### v2

@@ V Methods
We employ the approach of Krioukov et al. [52, 53], which allows one to generate random scale-free graphs $G(V,E)$ with controllable parameters, including the power-law degree distribution exponent $\gamma$. The network nodes are embedded in hyperbolic space $\mathbb{H}^{d}$ (with $d\geq 2$), where each node is assigned one hyperbolic radial coordinate $r$ and d1d-1 angular coordinates $\boldsymbol{\theta}^{\mathrm{emb}}=(\theta^{\mathrm{emb}}_{1},\theta^{\mathrm{emb}}_{2},\dots,\theta^{\mathrm{emb}}_{d-1})$, reflecting an effective geometric structure. For simplicity, we consider stimulus spaces that define an $d^{\mathrm{stim}}$-dimensional orientation (sphere $\mathbb{S}^{d^{\mathrm{stim}}}$). A stimulus is defined as a scalar field over $\mathbb{S}^{d^{\mathrm{stim}}}$, for example, exhibiting a cosine-like profile in the stimulus coordinates $\boldsymbol{\theta}^{\mathrm{stim}}=(\theta_{1}^{\mathrm{stim}},\theta_{2}^{\mathrm{stim}},\dots,\theta_{d^{\mathrm{stim}}}^{\mathrm{stim}})$. In both rate-based and spiking models of neural activity (see the next subsection), the $d^{\mathrm{stim}}$-dimensional stimulus space is associated with the d1d-1 -dimensional embedding boundary via the mapping $\boldsymbol{\theta}^{\mathrm{emb}}\mapsto\boldsymbol{\theta}^{\mathrm{stim}}$, under the dimensionality-matching condition $d^{\mathrm{stim}}=d-1$.
@@
where $f^{\mathrm{stim}}_{i}(t)$ is the value of the stimulus field in the receptive coordinates of neuron $i$, and $\xi_{i}(t)\sim\mathcal{N}(0,1)$ is Gaussian white noise. The parameters $g^{\mathrm{stim}}=5$, $g^{e,\mathrm{noise}}=1$, and $g^{i,\mathrm{noise}}=4$ scale the contributions of stimulus and noise to excitatory and inhibitory neurons, respectively.
As a stimulus field for detecting receptive fields in the spiking model, we use a narrow profile that smoothly decays toward the periphery, defined as $I^{\mathrm{stim}}(\boldsymbol{\theta})=\exp\left(-\|\boldsymbol{\theta}_{\mathrm{stim}}-\boldsymbol{\theta}\|/w_{\mathrm{stim}}\right)$, where $\boldsymbol{\theta}_{\mathrm{stim}}$ denotes the current peak location on the stimulus sphere, and $w_{\mathrm{stim}}$ controls the spatial width of the stimulus.
@@
where $S\in\mathbb{R}^{N\times N}$ is the internal connectivity matrix that determines how the nodes influence each other. Excitatory connections are defined according to the adjacency matrix of a scale-free graph generated and are scaled by excitatory conductance $g^{e}$. All remaining entries in the matrix, that is, those that do not correspond to excitatory connections, represent inhibitory influences and are randomly sampled from a uniform distribution in the interval [1,0][-1,0], then scaled by inhibitory conductance $g^{i}$. In this way, the matrix $S$ simultaneously encodes structured excitatory input and homogeneous random inhibitory connectivity. Synaptic depression modifies transmission through the probability of release $P_{\mathrm{rel}}(t)$  [70], which evolves as
@@
where $\tau_{\mathrm{depr}}=500$ ms is the recovery time constant, $P_{0}=1$ is the baseline release probability, and $f=0.8$ is the depression factor. The depression factor $f=0.8$ describes how much the synapses depress after each spike, effectively modifying the release probability via $P_{\mathrm{rel}}\rightarrow fP_{\mathrm{rel}}$. The firing rate is determined by a simple threshold function $r(t)=[I(t)]_{+},$ where $\cdot$$_{+}$ denotes rectification, corresponding to taking the non-negative part: $x]_{+}=\max(x,0)$.
@@
A full description of the methodology is provided in  [66]. The CA1 hippocampal activity was recorded in rats navigating a linear track with a total length of approximately $5$ meters, composed of ten $\sim 50$ cm sections. These segments were arranged into a flexible U-shaped track. The two arms of the track moved independently by remotely operated stepper motors. The displacement varied pseudo-randomly across trials, allowing to test a large set of configurations. To control distal visual cues, experiments were conducted in complete darkness. To support navigation, the track surface was coated with a scentless, washable, nontoxic glow-in-the-dark paint, which provided a faint green glow. A large screen blocked the view of one glowing arm from the other, ensuring the rat could only see the arm it was traversing. Stepper motors operated quietly, and the animals showed no overt behavioral responses to the associated noise or vibrations during track repositioning. Each animal was recorded for $7$ days on the track, performing on average about $40$ laps during each run session. The average traversal time ($\sim 34$ s) did not significantly differ from pre-training performance on the static track. More details can be found in [66].
@@
All analyses excluded epochs when animals ran at speeds <<2 cm/s. To compare spiking activity in planar and linear reference frames, we used Cartesian 2D coordinates and then linearized the trajectory such that the $x$ coordinate represented distance from one of the food wells, ignoring displacements orthogonal to the track direction. Receptive fields were estimated from single-unit spiking activity along the linear track. For each cell, spike positions were binned along the $x$-coordinate and converted into a spatial firing-rate profile by normalizing spike counts with the occupancy, i.e. the time spent by the animal in each spatial position. The resulting firing-rate profiles were smoothed with Gaussian filtering to reduce noise while preserving the structure of place fields. The reception fields were then defined as contiguous regions where the smoothed firing rate exceeded a fixed threshold of 2 Hz above baseline activity. For consistency of analysis, only a single receptive field was retained per neuron: in cases where multiple fields were detected, all but the largest contiguous field containing the peak firing location were discarded. Data analysis was performed in Python. To quantify the curvature of the receptive field distribution, we followed the exponential model of place-field sizes described by Eq. (1):
@@
where $s$ denotes the field size, $\xi$ is the global curvature parameter, and $s_{\min},s_{\max}$ are the minimal and maximal observed field sizes. To estimate the local curvature $\xi_{\mathrm{loc}}$ along the track, we applied this formulation within sliding windows (40 cm) on the spatial axis $x$. In each window, receptive field sizes $s$ with centers located inside the window were extracted, and a maximum likelihood estimate of $\xi$ was computed using the truncated exponential model above. This procedure yielded a spatially resolved profile $\xi_{\mathrm{loc}}(x)$, capturing local variations in field-size scaling. The resulting $\xi_{\mathrm{loc}}(x)$ values were subsequently smoothed with a Gaussian kernel to suppress noise while preserving spatial trends.